# Structure-Dynamics Correlation and Its Link to Fragility and Dynamic Heterogeneity

### v2

@@ IV.3 Dynamic heterogeneity and structure-dynamics correlation
Figure 11: Top panels: Partial radial distribution function, $g_{AA}(r)$, and interparticle potential, $U_{AA}(r)$, plotted as functions of $r$ for three sets of systems: (a)LJ and (b)WCA, each examined at four different densities, and (c) modified LJ $(q,p)$ system. Bottom panels: Spearman rank correlation,$C_{R}((\beta\phi)^{-1},X)$, between particle mobility, $\mu$, and coarse-grained structural parameters $X$—where $X$ represents SOP, $(\beta\phi)^{-1}$ (circles), two-body entropy, $S_{2}$ (squares), and local energy, $e_{\text{loc}}$ (triangles)—plotted as a function of coarse-graining length, $L$, for the same systems shown in top panels. The colour coding in the bottom panels matches that of the top panels.
@@
In the literature, dynamical heterogeneity is characterised using several complementary approaches. In this study, we employ the logarithm of the distribution of single-particle displacements at the $\alpha$- relaxation time, expressed as, $P(log_{10}(\delta r(t=\tau_{\alpha}))=ln(10)4\pi\delta r^{3}G_{s}(\delta r,t)$ which is related to the self-part of the van Hove correlation function, $G_{s}(\delta r,t)$ Flenner2005_pre . In addition, we analyse two commonly used dynamical measures: the four-point dynamic susceptibility $\chi_{4}$,Flenner2014_prl , and the non-Gaussian parameter $\alpha_{2}$ Rahman1964_review . Systems exhibiting stronger dynamical heterogeneity typically show a pronounced long tail in, $G_{s}(\delta r,t)$ and a double peak structure in $P(log_{10}(\delta r(t=\tau_{\alpha}))$, and both the $\chi_{4}$ and the $\alpha_{2}$ are found to have higher values Barrat2010_soft ; Lacevic2003_jcp ; Flenner2014_prl ; Kob1997_prl . To further probe dynamic variability, we also examine the distribution of particle mobility $\mu$ obtained from isoconfigurational ensemble simulations. The mobility distribution, $P(\mu)$ is not identical but similar to the $4\pi\delta r^{2}G_{s}(\delta r,t)$. The former represents a averaged measure of motion over many velocity realizations and thus suppresses some of the heterogeneity captured in the latter. Along with the distribution of $\mu$, we also calculate the distribution of the $\alpha$- relaxation time from the single particle overlap function, where the overlap function is averaged over the different velocity realizations. In Fig.8 we plot the functions for the LJ system, in Fig.9 we plot the same for the WCA system and in Fig.10 we plot them for the modified LJ $(q,p)$ system.
@@
Our analysis reveals that different measures of dynamic heterogeneity do not necessarily track each other. Interestingly, the systems exhibiting the maximum structure–dynamics correlation, LJ at $\rho=1.1$ and $(8,5)$ are both characterized by broader mobility distributions (Fig.8(a) and Fig.10(a)), broader distribution of $P(\tau_{\alpha})$ (Fig.14(a) and Fig.14(e) in the Appendix IV), prominent two-peaked structure of $P(log_{10}(\delta r(t=\tau_{\alpha}))$ (Fig.8(b) and Fig.10(b)) and a higher $\alpha_{2}$ value (Fig.8(d) and Fig.10(d)). However both these systems show low values of $\chi_{4}$ (Fig.8(c) and Fig.10(c)). Interestingly, the LJ at $\rho=1.1$ within its set is the least fragile system but the $(8,5)$ within its set is the most fragile system. For the LJ system, except for $\chi_{4}$ no other functions follow fragility, and it is just the reverse for the modified LJ $(q,p)$ system where $\chi_{4}$ does not follow fragility, but all the other functions do. For the WCA system where the change in fragility is the widest, the distribution of mobility $P(\mu)$ (Fig.9(a)) and the $\alpha$ relaxation time $P(\tau_{\alpha})$ (Fig.14(c) in the Appendix) are quite similar for all the densities however the other functions, $P(log_{10}(\delta r(t=\tau_{\alpha}))$, $\alpha_{2}$ and $\chi_{4}$ follow fragility.
@@ V Conclusion
The main objective of this study is to investigate the relationship between structure–dynamics correlation and fragility in supercooled liquids. Fragility is often associated with enhanced dynamic heterogeneity Sastry2001_nature ; Ediger2000_Physchem . Thus, if structural features significantly influence dynamics, one would expect systems with higher fragility to also exhibit stronger structure–dynamics correlation. In a previous study involving some of us, it was shown that the slope of the macroscopic energy barrier, obtained from average dynamics as a function of the structural order parameter (SOP), correlates with fragility Nandi2021_prl . A more recent study demonstrated that the slope of the microscopic energy barrier, when plotted against a machine-learned softness parameter, also increases with fragility Tah2022_jcp . Physically, the slope of the energy barrier versus SOP quantifies how sensitively the barrier, and thus the dynamics, responds to structural changes: a higher slope indicates that even small variations in SOP lead to significant changes in the barrier, reflecting stronger structure–dynamics coupling.
@@
Our SOP is observed to vary linearly with temperature, which allows the relaxation time to be fitted in terms of the SOP using a Vogel-Fulcher-Tammann (VFT) formGarciaColin1989_vft . The fragility extracted from this SOP-based VFT expression is consistent with the temperature-based fragility. However, a key insight from this analysis is that the slope of the energy barrier as a function of SOP does not necessarily scale with fragility, as often assumed Nandi2017_prl ; Tah2022_jcp . We also find that the barrier contains a temperature/SOP independent term, which sets the high-temperature limit of the activation energy and affects the slope when unaccounted for. At the macroscopic level, this scaled slope does align with fragility trends. However, at the microscopic level, the relationship is not universal. Notably, an independent and complementary analysis based on isoconfigurational ensemble simulations reveals that the structure–dynamics correlation correlates well with the microscopic scaled slope, reinforcing the idea that high structure–dynamics correlation does not necessarily coincide with high fragility.
@@
This raises the question of whether enthalpy dominance is intrinsically linked to broad mobility distributions and a form of heterogeneity. Interestingly, both the LJ at $\rho=1.1$ and the $(8,5)$ systems lie close to the spinodal line, a regime known to amplify dynamic heterogeneity due to mechanical instability and growing fluctuations Testard2014_jcp . Although the mechanisms underlying heterogeneity near the spinodal and in supercooled liquids differ, both are characterized by spatially heterogeneous mobility. In systems like $(8,5)$, high fragility and spinodal proximity may act in tandem to produce pronounced dynamical heterogeneity.
In conclusion, our findings reveal that fragility and structure–dynamics correlation do not always correlate. While fragility has often been associated with increased dynamic heterogeneity, our study shows that this relationship is more nuanced. Importantly, dynamic heterogeneity must be evaluated using a broader set of metrics beyond the commonly used four-point susceptibility, which captures spatially correlated motion. We find that the breadth of mobility distributions, such as those obtained from isoconfigurational ensemble simulations, is often a more sensitive indicator of the structural influence on dynamics than $\chi_{4}$. This perspective broadens our understanding of the glass transition and underscores the necessity of employing multiple, complementary measures when exploring the complex interplay between structure, dynamics, and fragility in supercooled liquids.
@@
We work with the absolute value of the caging potential, and the SOP which is the inverse of the depth of the caging potential, is given by,
[…]
@@
[…]
We present the temperature dependence of the SOP for all three sets of systems studied (LJ, WCA, and modified LJ $(q,p)$ system). In the right panels of Fig.2 we show that both $1/(\beta\Phi)$ and $<1/(\beta\phi)>$ exhibit a linear dependence on temperature across all systems. This observation is consistent with earlier studies Nandi2021_prl ; Sharma2022_pre . Although the macroscopic and microscopic SOP have the same functional forms, it is important to note that since the SOP is a nonlinear function of the radial distribution function, the average of the microscopic SOP is quantitatively not the same as the macroscopic SOP(Eq.3). Although $\langle g^{i}_{uv}(r)\rangle=g_{uv}(r)$, $\langle C^{i}_{uv}(r)g^{i}_{uv}(r)\rangle\neq C_{uv}(r)g_{uv}(r)$, i.e., the average of a product is not equal to the product of averages. As a result, the macroscopic SOP $(\beta\Phi)^{-1}$ and microscopic average SOP $\langle(\beta\phi)^{-1}\rangle$ have different values but interestingly similar trend as seen in right panels of Fig.2.
@@
where $\Delta E(\beta\phi)$ is interpreted as the microscopic activation energy required for rearrangement and depends on the local structure via $(\beta\phi)$. This is shown in Fig.12(b), which demonstrates the increasing structural control over dynamics at lower temperatures. This method thus establishes a direct quantitative link between local structural features and the energy barrier for particle rearrangements.
@@
To support the fragility analysis discussed in the main text, we present additional plots here that clarify the underlying trends for all three sets of systems: LJ, WCA, and modified LJ $(q,p)$ system. In the left panels of Fig.13, we show $\Delta E^{ma}$ plotted against $1/T$. These plots reflect how activation energy varies with temperature and how its slope correlates with the system’s fragility. As mentioned in the main text, for LJ and WCA systems, increasing density results in higher $T_{VFT}$ and fragility, accompanied by an increase in the slope of $\Delta E^{ma}$ vs $1/T$. In contrast, among the modified LJ $(q,p)$ system, $(8,5)$ exhibits the highest fragility but the lowest $T_{VFT}$, resulting in the smallest slope. In the right panels of Fig.13, we show the corresponding plots of $1/\Delta E^{ma}$ versus $1/T$. The slopes of these linear fits yield the fragility, $K_{\mathrm{VFT}}$, which exactly matches the fragility values obtained from the VFT fit of the relaxation time using Eq.11. The extracted values are listed in Table 1.
[…]
@@
Figure 15: Left panels: Spearman rank correlation,$C_{R}((\beta\phi)^{-1},\mu)$, between the SOP, $(\beta\phi)^{-1}$, and mobility,$\mu$, computed at $L=1$ for three different sets of systems:(a)LJ and (c) WCA, each examined at four different densities, and (e) modified LJ $(q,p)$ system. This figure complements the main results shown in Fig.7, and demonstrates that the observed density-dependent behavior is not sensitive to the choice of $L$. Right panels: Spearmann rank correlation,$C_{R}((\beta\phi)^{-1},X)$, between the SOP, $(\beta\phi)^{-1}$, and the dynamic quantity $X$, where $X$ represents either the instantaneous mobility $\mu$ evaluated at the $\alpha$-relaxation time (closed symbols) or the particle-level $\alpha$-relaxation times averaged over 200 isoconfigurational ensembles (open symbols) for the same systems shown in left panels. The colour coding in the right panels matches that of the left panels.

### v4

@@ IV.3 Dynamic heterogeneity and structure-dynamics correlation
[…]
@@
In the literature, dynamical heterogeneity is characterised using several complementary approaches. In this study, we employ the logarithm of the distribution of single-particle displacements at the $\alpha$- relaxation time, expressed as, $P(log_{10}(\delta r(t=\tau_{\alpha}))=ln(10)4\pi\delta r^{3}G_{s}(\delta r,t)$ which is related to the self-part of the van Hove correlation function, $G_{s}(\delta r,t)$ Flenner2005_pre . In addition, we analyse two commonly used dynamical measures: the four-point dynamic susceptibility $\chi_{4}$,Flenner2014_prl , and the non-Gaussian parameter $\alpha_{2}$ Rahman1964_review . Systems exhibiting stronger dynamical heterogeneity typically show a pronounced long tail in, $G_{s}(\delta r,t)$ and a double peak structure in $P(log_{10}(\delta r(t=\tau_{\alpha}))$, and both the $\chi_{4}$ and the $\alpha_{2}$ are found to have higher values Barrat2010_soft ; Lacevic2003_jcp ; Flenner2014_prl ; Kob1997_prl . To further probe dynamic variability, we also examine the distribution of particle mobility $\mu$ obtained from isoconfigurational ensemble simulations as $\mu^{j}(t)=\frac{1}{N_{\text{IC}}}\sum_{i=1}^{N_{\text{IC}}}\left|\mathbf{r}_{i}^{j}(t)-\mathbf{r}_{i}^{j}(0)\right|$, where $\mu^{j}(t)$ is the average displacement of the $j^{\text{th}}$ particle over $N_{\text{IC}}$ isoconfigurational trajectories at time $t$. The mobility distribution, $P(\mu)$ is not identical but similar to the $4\pi\delta r^{2}G_{s}(\delta r,t)$. The former represents a averaged measure of motion over many velocity realizations and thus suppresses some of the heterogeneity captured in the latter. Along with the distribution of $\mu$, we also calculate the distribution of the $\alpha$- relaxation time from the single particle overlap function, where the overlap function is averaged over the different velocity realizations. In Fig.8 we plot the functions for the LJ system, in Fig.9 we plot the same for the WCA system and in Fig.10 we plot them for the modified LJ $(q,p)$ system.
@@
Our analysis reveals that different measures of dynamic heterogeneity do not necessarily track each other. Interestingly, the systems exhibiting the maximum structure–dynamics correlation, LJ at $\rho=1.1$ and $(8,5)$ are both characterized by broader mobility distributions (Fig.8(a) and Fig.10(a)), broader distribution of $P(\tau_{\alpha})$ (Fig.14(a) and Fig.14(e) in the Appendix IV), prominent two-peaked structure of $P(log_{10}(\delta r(t=\tau_{\alpha}))$ (Fig.8(b) and Fig.10(b)) and a higher $\alpha_{2}$ value (Fig.8(d) and Fig.10(d)). However both these systems show low values of $\chi_{4}$ (Fig.8(c) and Fig.10(c)). Interestingly, the LJ at $\rho=1.1$ within its set is the least fragile system but the $(8,5)$ within its set is the most fragile system. For the LJ system, except for $\chi_{4}$, no other functions follow fragility, and it is just the reverse for the modified LJ $(q,p)$ system, where $\chi_{4}$ does not follow fragility, but all the other functions do. For the WCA system where the change in fragility is the widest, the distribution of mobility $P(\mu)$ (Fig.9(a)) and the $\alpha$ relaxation time $P(\tau_{\alpha})$ (Fig.14(c) in the Appendix) are quite similar for all the densities however the other functions, $P(log_{10}(\delta r(t=\tau_{\alpha}))$, $\alpha_{2}$ and $\chi_{4}$ follow fragility.
@@
Our results indicate that while fragility generally correlates with $\chi_{4}$, which captures the spatial extent of correlated regions, it does not always align with other measures of dynamic heterogeneity that reflect the width of the distribution of individual particle mobility. In contrast, the structure–dynamics correlation is robustly linked to the width of the mobility distribution and does not depend on the spatial correlation length and, consequently, the fragility.
@@ V Conclusion
The main objective of this study is to investigate the relationship between structure–dynamics correlation and fragility in supercooled liquids. Fragility is often associated with enhanced dynamic heterogeneity Sastry2001_nature ; Ediger2000_Physchem . Thus, if structural features significantly influence dynamics, one would expect systems with higher fragility to also exhibit stronger structure–dynamics correlation. In a previous study involving some of us, it was shown that the slope of the macroscopic free energy barrier, obtained from average dynamics as a function of the structural order parameter (SOP), correlates with fragility Nandi2021_prl . A more recent study demonstrated that the slope of the microscopic free energy barrier, when plotted against a machine-learned softness parameter, also increases with fragility Tah2022_jcp . Physically, the slope of the free energy barrier versus SOP quantifies how sensitively the barrier, and thus the dynamics, responds to structural changes: a higher slope indicates that even small variations in SOP lead to significant changes in the barrier, reflecting stronger structure–dynamics coupling.
@@
Our SOP is observed to vary linearly with temperature, which allows the relaxation time to be fitted in terms of the SOP using a Vogel-Fulcher-Tammann (VFT) formGarciaColin1989_vft . The fragility extracted from this SOP-based VFT expression is consistent with the temperature-based fragility. However, a key insight from this analysis is that the slope of the free energy barrier as a function of SOP does not necessarily scale with fragility, as often assumed Nandi2017_prl ; Tah2022_jcp . We also find that the barrier contains a temperature/SOP independent term, which sets the high-temperature limit of the activation energy and affects the slope when unaccounted for. At the macroscopic level, this scaled slope does align with fragility trends. However, at the microscopic level, the relationship is not universal. Notably, an independent and complementary analysis based on isoconfigurational ensemble simulations reveals that the structure–dynamics correlation correlates well with the microscopic scaled slope, reinforcing the idea that high structure–dynamics correlation does not necessarily coincide with high fragility.
@@
This raises the question of whether enthalpy dominance is intrinsically linked to broad mobility distributions and a form of heterogeneity. Interestingly, both the LJ at $\rho=1.1$ and the $(8,5)$ systems lie close to the spinodal line Sastry2000_prl_press ; Sengupta2011_jcp (see Appendix V), a regime known to amplify dynamic heterogeneity due to mechanical instability and growing fluctuations Sastry2000_prl_press . Although the mechanisms underlying heterogeneity near the spinodal and in supercooled liquids differ, both are characterised by spatially heterogeneous mobility. In systems like $(8,5)$, high fragility and spinodal proximity may act in tandem to produce pronounced dynamical heterogeneity.
In conclusion, our findings reveal that fragility and structure–dynamics correlation do not always correlate. We would like to clarify an important distinction regarding different measures of dynamic heterogeneity. The peak of the four-point susceptibility, $\chi_{4}$, captures the spatial extent of correlated regions in the dynamics and is commonly used to quantify heterogeneity that grows with fragility. In contrast, other indicators of dynamic heterogeneity, such as the distribution of particle mobility or the non-Gaussian parameter, reflect the contrast in individual particle mobilities, which can be substantial even when $\chi_{4}$ is small. Our results show that the structure–dynamics correlation is more strongly associated with this contrast in mobility than the spatial correlation length captured by $\chi_{4}$. Consequently, systems with high structure–dynamics correlation do not necessarily exhibit high fragility or large $\chi_{4}$ values. This distinction highlights that different aspects of dynamic heterogeneity need not vary together and that structure–dynamics correlation can be strong even in systems with low fragility.
@@
We work with the absolute value of the caging potential, and the SOP, which is the inverse of the depth of the caging potential, is given by,
$SOP=(\beta\phi^{i}_{u})^{-1}=1/\Big{(}\rho\int\textbf{dr}\sum_{v}x_{v}C_{uv}^{i}(r)\,g_{uv}^{i}(r)\Big{)}$ (22)
@@
$\begin{split}&\langle(\beta\phi^{i}_{u})^{-1}\rangle=\left\langle 1/\Big{(}\rho\int\textbf{dr}\,\sum_{v}x_{v}C^{i}_{uv}(r)g_{uv}^{i}(r)\Big{)}\right\rangle_{i}\\ &=1/\Big{(}\rho\int\textbf{dr}\,\sum_{v}\langle C^{i}_{uv}(r)g_{uv}^{i}(r)\rangle x_{v}\Big{)}\end{split}$ (23)
We present the temperature dependence of the SOP for all three sets of systems studied (LJ, WCA, and modified LJ $(q,p)$ system). In the right panels of Fig.2 we show that both $1/(\beta\Phi)$ and $1/(\beta\phi)>$ exhibit a linear dependence on temperature across all systems. This observation is consistent with earlier studies Nandi2021_prl ; Sharma2022_pre . Although the macroscopic and microscopic SOP have the same functional forms, it is important to note that since the SOP is a nonlinear function of the radial distribution function, the average of the microscopic SOP is quantitatively not the same as the macroscopic SOP(Eq.3). Although $\langle g^{i}_{uv}(r)\rangle=g_{uv}(r)$, $\langle C^{i}_{uv}(r)g^{i}_{uv}(r)\rangle\neq C_{uv}(r)g_{uv}(r)$, i.e., the average of a product is not equal to the product of averages. As a result, the macroscopic SOP $(\beta\Phi)^{-1}$ and microscopic average SOP $\langle(\beta\phi)^{-1}\rangle$ have different values but interestingly similar trend as seen in right panels of Fig.2.
@@
where $\Delta E(\beta\phi)$ is interpreted as the microscopic activation energy required for rearrangement and depends on the local structure via $(\beta\phi)$. This is shown in Fig.12(b), which demonstrates the increasing structural control over dynamics at lower temperatures. This method thus establishes a direct quantitative link between local structural features and the free energy barrier for particle rearrangements.
@@
The temperature dependence of the barrier $\Delta E^{ma}=T*ln(\tau_{\alpha}/\tau_{0})$ as predicted by the VFT expression (Eq.11) can be written as,
$\Delta E^{ma}/T=\frac{T_{VFT}}{K_{VFT}(T-T_{VFT})}$ (26)
This clearly shows that the slope of $\Delta E^{ma}$ is not directly proportional but inversely proportional to the fragility. For both the LJ and WCA systems, fragility increases with density, accompanied by a corresponding rise in $T_{VFT}$ as the systems at higher densities operate in higher temperature regimes. This trend explains why we observe an increase in the slope, $\frac{T_{VFT}^{2}}{K_{VFT}}$, with increasing fragility. For the modified LJ $(q,p)$ system the fragility is higher for the $(8,5)$ system which has the lowest value of $T_{VFT}$ and this is the reason even in the $\Delta E^{ma}$ vs $1/T$ plots the $(8,5)$ system has the slowest growth (left panels of Fig.13). Interestingly the slope of $1/\Delta E^{ma}$ vs $1/T$ is the fragility (right panels of Fig.13),
$1/\Delta E^{ma}=\frac{K_{VFT}}{T_{VFT}}-K_{VFT}.\frac{1}{T}$ (28)
The slopes of these linear fits yield the fragility, $K_{\mathrm{VFT}}$, which exactly matches the fragility values obtained from the VFT fit of the relaxation time using Eq.11. The extracted values are listed in Table 1.
Figure 13: Left panels: Variation of the macroscopic activation energy $\Delta E^{ma}$ with inverse temperature $1/T$ for three sets of systems: (a)LJ and (c)WCA, each examined at four different densities, and (e) modified LJ $(q,p)$ system. Right panels:The inverse activation energy, $1/\Delta E^{ma}$, is plotted against $1/T$ for the same systems. The slope of the resulting linear fits provides the fragility, $K_{\mathrm{VFT}}$ as shown in Eq.28. The corresponding values are listed in Table 1. The colour coding in the right panels matches that of the left panels.
@@
[…]
@@
Appendix V: Detection of spinoidal point - Pressure vs. density curve
Fig.16 shows the variation of inherent structure pressure, $P_{eIS}$, as a function of density, $\rho$, for the Lennard-Jones (LJ) system (black circles) and the (8,5) system (red squares). For each data point, $P_{eIS}$ was obtained by averaging more than 1000 independent configurations of the inherent structure to ensure statistical reliability. The plot reveals a distinct minimum in $P_{eIS}$, located at $\rho=1.095$ for the LJ system and at $\rho=1.16$ for the (8,5) system. These values are similar to those reported earlier Sastry2000_prl_press ; Sengupta2011_jcp , and these minima are indicative of the mechanical instability associated with the spinodal point in the energy landscape. In particular, the densities at which we perform our simulations, $\rho=1.1$ for the LJ system and $\rho=1.2$ for the (8,5) system, lie close to these spinodal densities, highlighting that the systems are being studied near the edge of stability.
Figure 16: The inherent structure pressure, $P_{\text{eIS}}$, is plotted as a function of density, $\rho$, for the LJ system (black circles) and the (8,5) system (red squares). The minimum of $P_{eIS}$ occurs at $\rho=1.095$ for the LJ system and at $\rho=1.16$ for the (8,5) system.